# Comparison of two Cooperative Maneuver Planning Approaches at a Real-World T-Junction

### v2

@@ 5.2 Graph-based Representation and Network
The graph based scene representation used in this study is based on our proposal in [3], which proved to be suited for learning a sensible mixed traffic capable RL policy. Thus, the current traffic scene observation is defined as , where denotes the set of vertices that the vehicles are mapped to, as depicted in Fig. 4. The nodes of vehicles that are in conflict and thus need to be coordinated are connected by directed edges from the set . Different kinds of interactions are encoded by means of edge types. The crossing edge type indicates a pair of vehicles that are located in front of the intersection driving on paths that intersect or merge on the intersection area. Leader-follower relations are encoded by an edge of type same lane pointing from the leader to the following vehicle. This edge direction induces a influence of the leading vehicle on the follower, but not vice versa.
While CAVs share their desired route with the cooperative planner, the turn direction of human drivers is unknown. Therefore, the cooperative planner has to consider all possible conflicts with HDVs when coordinating CAVs. The graph-based scene representation inherently supports this by adding edges to an HDV for all conflicts that cannot be ruled out reliably. In the present work, we match the vehicle’s pose to lanes in a high definition (HD) map and enumerate viable continue paths. This assignment allows an angle between the vehicle heading the lane of up to at the matched point. When crossing the intersection, a vehicle’s lane matches will thus reduce to the actual destination eventually, which then allows to plan the maneuver with less excess edges. The delay of this decision might be reduced further by employing a prediction algorithm, such as [33], which is out of scope for this work, though.
In addition to the semantic structure encoded in the graph, both vertices and edges are augmented by input feature sets. The vertex input features contain the vehicle’s longitudinal position along its lane, the driving speed, and measured acceleration. Additionally, a binary flag indicates whether the corresponding vehicle is a CAV to be coordinated or a HDV. The edge features are composed of a distance measure, heading-relative bearing, and the priority according to the map. The latter is required to enable the cooperative planner to correctly anticipate the behavior of HDVs that obey to the legacy priority rules.
The RL policy shall derive a joint action for all CAVs in the scene composed of longitudinal acceleration commands in continuous space for each vehicle. We employ the TD3 [34] actor-critic RL algorithm to train a GNN for the cooperative planning task. For conciseness, we present the actor network architecture in Fig. 5, while the critic’s architecture deviates only slightly on the output side. The input features for both vertices and edges are first encoded by MLPs with shared weights, respectively. For message passing in the GNN, we use modified relational graph convolutional network (RGCN) layers [35] as well as graph attention (GAT) layers [36]. The RGCN layers maintain independently learnable weight matrices per edge type and also incorporate the edge features in the vertex feature update. By introducing one intermediate GAT layer, the network may leverage the attention mechanism to focus on the most relevant interactions. A pure GAT network is not suited though, as the edge features are only used to compute the keys and queries, but not the values for the vertex feature update. Moreover, GAT does not consider the different edge types that are crucial for our graph-based representation.
@@
The resulting vertex features are finally mapped to a joint action by another MLP in the actor architecture. In contrast, the critic network aggregates the vertex feature vectors resulting from message passing to a single feature vector before decoding a Q-value estimate using an MLP. The graph-based scene representation and the GNN are implemented using the PyTorch Geometric API [37].
5.3 Reward Function Engineering
Apart from a suited network architecture, RL requires a well-designed reward function to learn a reasonable policy. For solving the cooperative planning problem through RL, the reward function is defined as
with describing the set of reward components and the corresponding weights. The only positive reward component and thus the main driver for learning a nontrivial solution is the velocity reward, which is proportional to the average driving speed clipped to the lane speed limit. A trivial solution that consequently avoids any collision is to stop all vehicles in the scene. This is not a valid solution for cooperative maneuver planning and thus prevented through the idle penalty. In mixed traffic, the training tends to converge to a very conservative behavior, which manifests in all CAVs being stopped far away from the intersection to let HDVs pass. This issue is addressed by the reluctance penalty, which is triggered for CAVs that drive very slowly without any lead vehicle. In order to facilitate learning of reasonable safety distances, the proximity reward penalizes situations in which CAVs get dangerously close. Ultimately, two vehicles colliding triggers the collision penalty and ends the training episode to avoid further positive rewards from being accumulated. The reward weights used throughout this study, at a policy frequency of , are given in Table 2. Details on the definition and tuning of the various reward components can be found in [24, 3].
Reward | velocity | idle | reluctance | proximity | collision |
Weight | 0.06 | 0.02 | 0.02 | 0.2 | 1.0 |
5.4 Derivation of Cooperative Maneuver
The maneuver representation introduced in Sec. 3 requires the cooperative maneuver to be specified over a sufficient horizon to facilitate local trajectory planning on the CAVs. This requirement for advance planning contrasts with the iterative nature of the RL policy, which requires immediate state feedback. In the present work, we build upon our approach from [26] to bridge this gap.
Based on the current traffic condition, a cooperative maneuver is derived by using a built-in simulator. After initializing the simulator based on the current state of the server-side EM, the RL policy is queried to predict the future evolution of the scene. We employ the same simulation environment that was used for training of the RL policy, based on Highway-env [38]. As depicted in Algorithm 1, the current simulator state is observed at and passed to the RL policy . While the simulated CAVs are moved according to the selected actions, the behavior of regular vehicles is predicted by the intelligent driver model (IDM, [39]), enhanced by the yielding logic from [25]. As the true destination of regular vehicles is not known to the cooperative planner, a worst-case route estimate is used. Thereby, all potential conflict points are considered while planning the cooperative maneuver.
The simulation run is terminated when all vehicles have reached their destination, e.g., the desired side of the intersection, or optionally after a timeout. Typically, the overall planning response time remains below , however, it might occasionally exceed in very crowded scenarios (more than ten vehicles present). In those cases, aborting the simulation by timeout is a viable measure to limit the response time. Because it is not ensured that every vehicle is guided to the end of the intersection area, cyclic replanning is required. In the current work, the cooperative maneuver is updated on a period, as it was shown in [26] that even slight deviations in the predicted behavior can lead to maneuvers becoming infeasible. Especially with HDVs being present, cyclic replanning becomes imperative to ensure drivable maneuvers.
@@ 5.4 Derivation of Cooperative Maneuver
For each recorded simulated trajectory, waypoint candidates are created at the entry and exit of the intersection conflict area. The lower and upper bounds for traversing these waypoints is obtained from the simulation. Thereby, the entry waypoint determines the earliest possible time of entering the intersection. Analogously, the waypoint at the intersection entry specifies when the receiving CAV shall clear the conflict area at the latest.
6 SIMULATIVE EXPERIMENTS
@@
We integrated the proposed cooperative planning approaches into an extended simulation and evaluation framework for fully automated and mixed traffic. The cooperative planners are evaluated and assessed against two baselines. Our primary baseline are non-cooperative (NC) CAVs that leverage shared perception data, but no routing information and adhere to priority rules. This baseline employs the same underlying trajectory planning as is used for the cooperative approaches. Thus, it constitutes a rather strong baseline, which facilitates a sharp evaluation of the cooperative aspect in isolation. In addition, we present the metrics for pure HDV traffic, which is simulated using a learned behavior model. As such, it resembles human driving behavior more closely and exhibits a human-typical agile behavior.
6.1 Simulation Setup
@@
The simulation framework that we used for evaluation of the planning methods builds upon the DeepSIL simulator introduced in [4]. It is based on the ROS2 software distribution [40] and can simulate HDVs as well as CAVs. We extended the framework to support multiple CAVs and cooperative maneuvers. The simulation has a base time step of and runs in real-time, meaning that processing delays from the cooperative maneuver planners directly affect the results. When simulating more than four CAVs, the simulation is slowed down slightly to still be able to run all vehicle trajectory planners in parallel.
@@
Human-driven vehicles are predicted using the approach in [4]. Every , the multiple trajectory prediction network is evaluated on the current traffic environment. In each cycle, the most progressing, non-conflicting prediction is used to update the vehicle position and velocity. The prediction is converted into a longitudinal action and the internal vehicle simulation is stepped forward along the centerline of the route. The intersection handling was adapted to avoid collisions even in edge cases appearing in a large number of simulations. Also, an occlusion simulation was added, i.e. HDVs required to yield assume a blocked intersection and approach slowly until they have clear vision of the conflicting routes. The simulated HDVs do not know the routes of the other vehicles and therefore need to assume the most conflicting route.
@@
Connected automated vehicles are simulated using a single-track model with friction and dead time, resembling one of our real-world CAVs. The inputs to this model are calculated by the trajectory planner and controller from [5], which is part of this CAV’s software and therefore provides a realistic CAV behavior. Like HDVs, CAVs do not know the route of other vehicles, even of other CAVs, and assume the most conflicting route. However, the simulated CAVs do not need to approach the intersection slowly due to occlusion, because they receive the intersection EM as in [1].
6.2 Scenario Definition
@@
We define all simulated scenarios at an exemplary intersection in Ulm-Lehr, Germany, as shown in Fig. 6, where also our pilot site (cf. Sec. 7.1) is located. It is a T-junction with a bending main road and a side road. The speed limit is () on the eastern and western part and () on the northern part.
@@
Within this setting, we created a total set of 480 scenarios that serve as an initial configuration to the simulation. In each scenario, a number of vehicles is spawned at the locations shown in Fig. 6, where each vehicle has an initial speed of and a specific route, i.e., turn direction. The first 280 scenarios are an enumeration of all combinations of source and target directions with up to two vehicles from each direction and in which there is at least one pair of conflicting vehicles. The remaining 200 scenarios are randomly sampled with up to four vehicles from each direction and a random route, limited to eight vehicles total due to computational constraints of the simulation.
@@
The scenario set was simulated both in fully automated traffic and mixed traffic. When simulating mixed traffic, the vehicles are randomly divided into HDVs and CAVs, where for an odd number of vehicles, the last one is chosen randomly. Out of the 480 total scenarios, less than of the simulations were not successful and resulted in a timeout or collision due to remaining issues in the simulation and trajectory planning frameworks (five scenarios in fully automated traffic, 14 scenarios in mixed traffic). Those scenarios were excluded from the evaluation in all maneuver planning configurations.
@@
To ensure a fair comparison, all trajectories have been clipped to common evaluation intervals, as illustrated in Fig. 6. The following evaluations refer to trajectory segments that begin in front of the intersection entry ( in Fig. 6) and end behind the intersection ( in Fig 6). These intervals cover the relevant part of the intersection approach during which the CAVs synchronize onto their assigned slot for crossing the intersection.
6.3 Quantitative Results
@@
Figure 7 depicts the results in fully automated traffic. Both planners are able to realize cooperative maneuvers in the vast majority of scenarios, as can be seen from Fig. LABEL:sub@sfig:all_cav_deviating. A scenario run is considered cooperative if the vehicle crossing order deviates from the baseline crossing order. For comparability, the following analyses are performed over all scenarios in which at least one cooperative planner suggested a deviating crossing order. In the cooperative scenarios, the traffic flow especially on the minor road is improved significantly, while the optimization-based planner provides an additional benefit over the RL-based one. Figure LABEL:sub@sfig:all_cav_relative_duration_road_type depicts the distribution of the duration required by the vehicles to complete the maneuver relative to non-cooperative CAVs. Almost all vehicles on the minor road benefit of a reduced delay due to interaction. Some vehicles on the major road take slightly longer to complete the maneuver in the cooperative case. This is expected since the vehicles on the major road waive their unconditional right of way when taking part in cooperative maneuvers. Overall, the cooperation leads to a reduction in median duration, which becomes even more pronounced when comparing to pure HDV traffic.
@@
Having a vehicle stop is particularly disadvantageous in terms of both energy efficiency and passenger comfort. As can be observed in Fig. LABEL:sub@sfig:all_cav_stopped_vehicles_road_type, both cooperative planners eliminate most of the stops required on the minor road when driving according to priority rules. Notably, the share of stopped vehicles on the major road does not increase but even decrease slightly, which can mainly be attributed to left-turning vehicles yielding to oncoming traffic. Pure HDV traffic exhibits much more stops because they cannot rely on shared perception data and thus have to approach the intersection more carefully.
@@
The cooperative planning approaches also raise improvement potential in mixed traffic efficiency, as illustrated by Fig. 8. Both the optimization-based planner and the RL planner suggest a cooperative maneuver in significantly fewer cases compared to fully automated traffic. It is worth noting that a share of HDVs renders many cooperative maneuvers infeasible because the priority relations towards HDVs have to be retained. The following metrics again refer to all scenarios in which at least one cooperative planner suggested a deviating crossing order.
@@
Figure LABEL:sub@sfig:mixed_relative_duration_road_type indicates that still many vehicles on the minor road benefit of a reduced delay due to the cooperative maneuver. It is important to note that the simulation setup only considers the evaluation of the initial vehicle configuration, but not continuous traffic. Unlike in real traffic, the vehicles on the minor road will be able to cross the intersection fluently after the prioritized traffic has left the intersection. This circumstance leads to a favoritism of the baseline maneuvers, which might have taken much longer in real continuous traffic. Nonetheless, a reduction of stops can be observed in Fig. LABEL:sub@sfig:mixed_stopped_vehicles_road_type also for mixed traffic. Here, the optimization-based planner clearly outperforms the RL-based planner in terms of stops on the minor road. This can most likely be attributed to the training of the RL being performed in continuous traffic to facilitate learning of a robust policy. In the case of evaluating single scenarios, the learned efficiency estimation of the RL model might not be accurate.
7 REAL-WORLD EXPERIMENTS

### v5

@@ III-D2 Graph-based Representation and Network
The graph based scene representation used in this study is based on our proposal in [19], which proved to be suited for learning a sensible mixed traffic capable RL policy. Thus, the current traffic scene observation is defined as a directed graph, as illustrated in Fig. 3. The nodes of vehicles that are in conflict and thus need to be coordinated are connected by directed edges. Due to the inherently unknown turning intention of HDVs, like in the figure, the graph comprises additional edges for all conflicts that cannot be ruled out reliably. Different kinds of interactions are encoded by means of edge types. The crossing edge type indicates a pair of vehicles that are located in front of the intersection driving on paths that intersect or merge on the intersection area. Leader-follower relations are encoded by an edge of type same lane pointing from the leader to the following vehicle.
In addition to the semantic structure encoded in the graph, both vertices and edges are augmented by input feature sets. While the vertex input features contain vehicle-relevant measurements, the edge features are composed of pair-wise properties like distances.
@@
The first part of the network architecture is comprised of MLPs that encode the vertex input features and edge input features, respectively. At the core of the GNN, modified relational graph convolutional network layers [33] and graph attention layers [37] are used for message passing in an interleaved structure. In the actor network architecture, the resulting vertex features are mapped to a joint action by a final output MLP. In contrast, the critic network aggregates the vertex feature vectors resulting from message passing to a single feature vector before decoding a Q-value estimate using another MLP. The graph representation and the GNN are implemented using the PyTorch Geometric API [10].
III-E Prediction and Derivation of Cooperative Maneuvers
For the optimization-based planner, the selected priority assignments need to be converted to maneuver constraints as in Eq. (1) for each individual CAV. Each pair of conflicting CAVs shares a conflict zone, i.e., a segment along the road that may only be occupied by one of the vehicles at any time to guarantee freedom from collision, as explained above. The predicted timestamp at which the prioritized vehicle leaves and the yielding vehicle may enter the conflict zone is used to encode the space-time maneuver constraints. The start and end points of the conflict zone along the respective vehicle routes are used as waypoints for the constraints.
@@ III-E Prediction and Derivation of Cooperative Maneuvers
The RL-based cooperative planner derives cooperative maneuvers using a built-in simulator. After initializing the simulator based on the current state of the server-side EM, the RL policy is queried to predict the future evolution of the scene. For each recorded simulated trajectory, waypoint candidates are dynamically created at the entry and exit of the intersection conflict zone. Similar to the optimization-based planner, the lower and upper time interval bounds are obtained from the simulation. Thereby, the entry waypoint determines the earliest possible time of entering the intersection, while the waypoint at the intersection exit specifies when the receiving CAV shall clear the conflict area at the latest.
IV Simulation Approach
@@
We integrated the proposed cooperative planning approaches into an extended simulation and evaluation framework for fully automated and mixed traffic. The cooperative planners are evaluated and assessed against two baselines. Our primary baseline are non-cooperative (NC) CAVs that leverage shared perception data as in [4], but no routing information and adhere to priority rules. This baseline employs the same underlying trajectory planning as is used for the cooperative approaches. Thus, it constitutes a rather strong baseline, which facilitates a sharp evaluation of the cooperative aspect in isolation. In addition, we present the metrics for pure HDV traffic, which is simulated using a learned behavior model. As such, it resembles human driving behavior more closely and exhibits a human-typical agile behavior. The focus of this paper is on the comparison of the two cooperative planning approaches at one exemplary intersection. We refer the reader to previous publications for more extensive individual evaluations of the optimization-based [25] and the RL-based [19] cooperative planners, respectively.
IV-A Simulation Setup
@@
The simulation framework that we used for evaluation of the planning methods builds upon the DeepSIL simulator introduced in [36]. It is based on the ROS2 software distribution [24] and can simulate HDVs as well as CAVs. We extended the framework to support multiple CAVs and cooperative maneuvers. The simulation has a base time step of 50 ms and runs in wall-clock time on a single machine. This means that processing delays from the cooperative maneuver planners are considered and affect the results, while communication delays are not modeled. Due to limited computational resources on the machine, simulations with more than four CAVs are slowed down slightly to still be able to run all vehicle trajectory planners in parallel.
@@
Human-driven vehicles are predicted using the approach in [36]. Every 50 ms, the multiple trajectory prediction network is evaluated on the current traffic environment. In each cycle, the most progressing, non-conflicting prediction is used to update the vehicle position and velocity. The prediction is converted into a longitudinal action and the internal vehicle simulation is stepped forward along the centerline of the route. A simple occlusion simulation was added, i.e., HDVs on the subordinate road approach slowly until they reach a distance of 2 m to the intersection, at which a full view into the situation is assumed. Also, the simulated HDVs do not know the routes of the other vehicles and therefore need to assume the most conflicting route.
@@
Connected automated vehicles are simulated using a single-track model with friction and dead time, resembling one of our real-world CAVs. The inputs to this model are computed by the trajectory planner and controller from [32], which is part of this CAV’s software and therefore provides a realistic CAV behavior. Like HDVs, CAVs do not know the route of other vehicles, even of other CAVs, and assume the most conflicting route. However, the simulated CAVs do not need to approach the intersection slowly due to occlusion, because they receive the intersection EM as in [4].
The two approaches for cooperative maneuver planning are developed and integrated as alternative ROS2 software modules. For each simulation run, only one of the modules is activated (or none in the non-cooperative case), while the remaining simulation setup stays identical. They communicate with the simulated CAVs and server-side EM using the V2X communication protocols we proposed in [4, 26].
IV-B Scenario Definition
@@
We define all simulated scenarios at an exemplary intersection in Ulm-Lehr, Germany, as shown in Fig. 4, where also our pilot site (cf. Section VI-A) is located. It is a T-junction with a bending main road and a minor subordinate road. The speed limit is () on the eastern and western part and () on the northern part.
@@
Within this setting, we created a total set of 480 scenarios that serve as an initial configuration to the simulation. In each scenario, a number of vehicles is spawned at the locations shown in Fig. 4, where each vehicle has an initial speed of and a specific route, i.e., turn direction. The first 280 scenarios are an enumeration of all combinations of source and target directions with up to two vehicles from each direction and in which there is at least one pair of conflicting vehicles. The remaining 200 scenarios are randomly sampled with up to four vehicles from each direction and a random route, limited to eight vehicles total due to computational constraints.
@@
The scenario set was simulated both in fully automated traffic and mixed traffic. When simulating mixed traffic, the vehicles are randomly divided into 50 % HDVs and 50 % CAVs, where for an odd number of vehicles, the last one is chosen randomly. Out of the 480 total scenarios, less than 3 % of the simulations were not successful and resulted in a timeout or software crash (five scenarios in fully automated traffic, 14 scenarios in mixed traffic). This is due to remaining subtle issues in the prototype CAV software stack and unrelated to the maneuver planning. Those scenarios were excluded from the evaluation in all maneuver planning configurations.
V Simulation Evaluation Results
@@
The extensive simulation experiments were quantitatively evaluated in terms of efficiency metrics as well as criticality metrics, which is presented in the following section. To ensure a fair comparison, all trajectories have been clipped to common evaluation intervals, as illustrated in Fig. 4. The following evaluations refer to trajectory segments that begin 60 m in front of the intersection entry ( in Fig. 4) and end 15 m behind the intersection ( in Fig 4). These intervals cover the relevant part of the intersection approach during which the CAVs synchronize onto their assigned slot for crossing the intersection.
V-A Effectiveness and Efficiency Analysis
@@
Fig. 5 depicts the results in fully automated traffic. Both planners are able to realize cooperative maneuvers in the vast majority of scenarios, as can be seen from Fig. LABEL:sub@sfig:all_cav_deviating. For comparability, the analyses in this section are performed over all cooperative scenarios, that is, all scenarios in which at least one cooperative planner deviates from the baseline crossing order. For all other scenarios, the metrics do not significantly differ from the baseline. In the cooperative scenarios, the traffic flow especially on the subordinate road is improved significantly, while the optimization-based planner provides an additional benefit over the RL-based one. Fig. LABEL:sub@sfig:all_cav_relative_duration_road_type depicts the distribution of the duration required by the vehicles to complete the maneuver relative to non-cooperative CAVs. Per definition, NC traffic exhibits a relative duration of and is thus not shown in the plot. Almost all vehicles on the subordinate road benefit of a reduced delay due to interaction. Some vehicles on the major road take slightly longer to complete the maneuver in the cooperative case. This is expected since the vehicles on the major road waive their unconditional right of way when taking part in cooperative maneuvers. Overall, the cooperation leads to a reduction in median duration, which becomes even more pronounced when comparing to pure HDV traffic. The optimization-based planner achieves slightly higher efficiency due to planning with a lower safety distance, as discussed in Section V-B.
@@
Having a vehicle stop is particularly disadvantageous in terms of both energy efficiency and passenger comfort. As can be observed in Fig. LABEL:sub@sfig:all_cav_stopped_vehicles_road_type, both cooperative planners eliminate most of the stops required on the subordinate road when driving according to priority rules. Notably, the share of stopped vehicles on the major road does not increase but even decrease slightly, which can mainly be attributed to left-turning vehicles yielding to oncoming traffic. Pure HDV traffic exhibits much more stops because they cannot rely on shared perception data and thus have to approach the intersection more carefully. Meanwhile, the NC baseline only leads to a moderate decrease in stops, despite its huge perception advantage.
@@
In mixed traffic, both cooperative planning approaches increase the traffic efficiency, as illustrated by Fig. 6, although to a lesser extent than in fully automated traffic. It is worth noting that a 50 % share of HDVs renders many cooperative maneuvers infeasible because the priority relations towards HDVs have to be retained. Thus, achieving a notable gain in efficiency is much more challenging in mixed traffic, which is consistent with the individual planner evaluations [25, 19].
@@
Fig. LABEL:sub@sfig:mixed_relative_duration_road_type indicates that still many vehicles on the subordinate road benefit of a reduced delay due to the cooperative maneuver. It is important to note that the simulation setup does not resemble continuous traffic but only the drive off of the initial vehicle configuration. Unlike in real traffic, the vehicles on the subordinate road will be able to cross the intersection fluently after the prioritized traffic has left the intersection. This circumstance leads to a favoring of the baseline maneuvers, which might have taken much longer in real continuous traffic. Nonetheless, a reduction of stops can be observed in Fig. LABEL:sub@sfig:mixed_stopped_vehicles_road_type.
The optimization-based planner clearly outperforms the RL-based planner in terms of stops on the subordinate road in mixed traffic. This can most likely be attributed to the training of the RL being performed in continuous traffic to facilitate learning of a robust policy. In the evaluated case of single scenarios, the learned efficiency estimation of the RL model might not be accurate. Also, the optimization-based planner seems to favor traffic efficiency while compromising in terms of safety distance, which is examined in the following section.
V-B Criticality Analysis
The gain in traffic efficiency due to cooperative maneuver planning should not be realized at the expense of safety. Thus, we complement the performance analysis by a maneuver criticality analysis based on three metrics:
The two-dimensional time-to-collision (TTC) is a generalization of the TTC metric to intersection scenarios;
The second criticality metric is defined as the deceleration required to avoid a collision (DRAC);
Finally, we report the post encroachment times (PET) for the maneuvers.
We calculate the TTC using the implementation from [16]. For the DRAC metric, the post-processing takes each simulated time step and calculates the required deceleration of each vehicle to avoid collision in case all other vehicles would suddenly decelerate with . The PET between two conflicting vehicles is the duration between the first vehicle leaving and the second vehicle entering the shared conflict zone.
For each simulated scenario, we gather the most critical (i.e., lowest TTC and PET, highest DRAC) metrics across all vehicles and time steps. Then, we report the median value across all evaluated scenarios and the percentage of critical scenarios (). For this classification, we apply commonly accepted thresholds of for TTC, for DRAC (both taken from [14]), and for PET [28].
The criticality results for fully automated traffic and for mixed traffic are depicted in Table II and Table III, respectively. Both planner methods yield a higher share of critical scenarios compared to the baseline, which is expected as they are optimized for efficiency. The RL planner results are closer to the NC baseline, which is most likely due to the strong focus on HDV interoperability in the reward scheme and training procedure. Meanwhile, the optimization-based planner shows significantly increased criticality percentages for DRAC and PET, which goes along with its better performance in traffic efficiency metrics. This is due to the fact that, internally, the planner calculates geometrically accurate conflict zones and thus assumes a PET of 0 s to be sufficiently safe. Increasing this safety distance to allow for more reaction time will be considered in future works.
| Planner | ||||||
| Opt | ||||||
| RL | ||||||
| NC |
| Planner | ||||||
| Opt | ||||||
| RL | ||||||
| NC |
VI Real-World Experiments